# Conditioning of Banach Space Valued Gaussian Random Variables: An Approximation Approach Based on Martingales

### v1

@@ Abstract
In this paper we investigate the conditional distributions of two Banach space valued, jointly Gaussian random variables. These conditional distributions are again Gaussian and their means and covariances are determined by a general approximation scheme based upon a martingale idea. We then apply our general results to the case of Gaussian processes with continuous paths conditioned to partial observations of their paths.
@@ 1 Introduction
Gaussian processes for machine learning (GP4ML), see e.g. [36, 5, 24], are a standard and widespread non-parametric Bayesian technique that allows for both prediction and uncertainty quantification in a single approach. In addition, their Bayesian update rule also plays, an important role in other fields such as information-based complexity, see e.g. [44], probabilistic numerics and uncertainty quantification, see e.g. [42, 9, 20], and Bayesian methods for inverse problems, see e.g. [41].
In the simplest setup, GP4ML assume that a Gaussian process $X:=(X_{t})_{t\in T}$ is given, where $T$ is some index set. Moreover, it is assumed that both the mean function $m:T\to\mathbb{R}$ and the covariance function $k:T\times T\to\mathbb{R}$ defined by
@@
Here, $K_{t_{1},S}$, $K_{S,S}$, and $K_{S,t_{2}}$ denote covariance matrices with respect to the indicated indices and $K_{S,S}^{\dagger}$ is the Moore-Penrose inverse of the $n\times n$-matrix $K_{S,S}$. Here we note that in the GP4ML literature, it is often assumed that we have noisy observations of the form $Y+\varepsilon$, where $\varepsilon$ is a centered Gaussian vector of i.i.d. entries that is independent of $Y$. In this case, the formulas slightly change, see the discussion following (29).
@@
Usually, the Gaussian process $X$ is viewed as a mechanism to generate random functions, e.g. random continuous functions on a compact metric space $T$. In this case, the “conditional Gaussian process” of $X$ given the observations $Y=y$ is also viewed as such a mechanism. To make this mathematically rigorous, one needs to view $X$ as a $C(T)$-valued Gaussian random variable, where $C(T)$ denotes the Banach space of continuous $\mathbb{R}$-valued functions on $T$, see Sections 2 and 4 for these and the following notions. Moreover, the above $Y$ is an $\mathbb{R}^{n}$-valued Gaussian random variable such that $X$ and $Y$ are jointly Gaussian. Given their joint distribution $\mathrm{P}_{(X,Y)}$ on $C(T)\times\mathbb{R}^{n}$, one then seeks a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$, since $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is the conditional distribution of $X$ given the observation $Y=y$. Note that by definition, each $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a distribution on $C(T)$ and therefore we can also ensure that the “conditional Gaussian process” again produces continuous functions.
@@
The references mentioned at the beginning clearly indicate that for the conditioning problem $C(T)$ is by no means the only Banach space one is interested in. Therefore, let us now consider the general case. To this end, let $E$ and $F$ be Banach spaces, which due to technical simplicity are assumed to be separable. Moreover, let $X$ be an $E$-valued random variable and $Y$ be an $F$-valued random variable such that $X$ and $Y$ are jointly Gaussian. In the abstract setting one then seeks a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$. Again, $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a distribution on $E$, which equals the conditional distribution of $X$ given the observation $Y=y$. Compared to the $C(T)$-case described earlier, this generalization opens several additional possibilities. First, of course, we can consider Banach spaces of functions $E$ other than $C(T)$, e.g. spaces of differentiable functions. Second, we no longer need to assume that we can observe the random functions on some finite number of observational points. Instead, we could assume, e.g., that we observe derivatives of the random functions at some observational points, cf. [35]. More generally, $Y$ can, for example, be composed by any other finite set of bounded linear functionals as considered in e.g. [45], or even by an infinite set of such functions. In particular, for the $C(T)$-example above, the abstract setting makes it possible to consider arbitrary subsets $S\subset T$ for the observational points, which is important for PDE-related applications, see e.g. [35]. To abstract from this, let us assume that $Y$ is given by
@@
Although working with infinite dimensional $Y$ has clear conceptual advantages, there is also a drawback: Indeed, in many situations, it is practically impossible to fully observe or process infinite dimensional $Y$. In these cases, one often uses a sequence $(Y_{n})$ of observable, finite dimensional Gaussian random variables, for which the condition step is feasible. For example, in the $C(T)$-case with infinite $S\subset T$, one could choose an increasing sequence $(S_{n})$ of finite subsets $S_{n}\subset S$ whose union is dense in $S$ and set $Y_{n}:=(X_{s})_{s\in S_{n}}$. The hope of this approach is, of course, that $\mathrm{P}_{X|Y_{n}}$ provides a convergent approximation of $\mathrm{P}_{X|Y}$, e.g. in terms of the means and covariances, or as distributions.
@@
Both the fully infinite dimensional conditioning and its possible approximations by finite dimensional observations have various applications as described in e.g. [44, 9, 20, 35]. Here, we only mention Bayesian quadrature, Bayesian optimization, and Bayesian solvers for differential equations. In addition, they can help to better understand GP4ML and their limiting behavior for $n\to\infty$. For this reason, a better understanding of the regular conditional probabilities of $X$ given $Y$ for jointly Gaussian, Banach space valued random variables $X$ and $Y$ as well as of possible approximations is of high practical interest.
@@
Let us therefore have a look what is known so far. To begin with, in the case of finite dimensional observations, that is $F=\mathbb{R}^{n}$, [6, Corollary 3.10.3] shows that there is a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$, such that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian for all $y$. However, no formulas for the mean and covariance of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ are provided. Moreover, the observational model (5) for centered $X$ and with arbitrary $E$ and $F$ has been investigated in [43], where the authors show that the regular conditional probabilities are Gaussian measures, see their Theorem 3.11. Moreover, in the case $F=\mathbb{R}^{n}$, their Corollary 3.13 present explicit formulas for the conditional mean and covariance with the help of a so-called representing sequence of the covariance operator $\operatorname{cov}Y$. In addition, a closer look at their proofs reveals, that analogous formulas hold true for general, separable $F$. The latter formulas have been made explicit and generalized to non-centered $X$ in [45]. Moreover, both papers establish the linearity of the conditional mean mapping, and [45] also describes formulas for sequential updates and the case of $C(T)$-valued $X$. In particular, for $E=C(T)$ and $F=\mathbb{R}^{n}$, the update rules (3) and (4) are generalized to arbitrary, bounded linear observational operators $\mathsf{S}:C(T)\to\mathbb{R}^{n}$. Unfortunately, however, the mean and covariance formulas for infinite dimensional $F$ become less handy, as they require to have a representing sequence for $\operatorname{cov}Y$. Although such sequences as well as constructions for them exist, see e.g. [43, Lemma 3.5] or [40, Lemma 8.2.3], it seems fair to say that they are rarely at hand in situations in which $F$ is not a Hilbert space. As a consequence, it seems fair to say that the derived formulas have some limited practical value.
@@
The case of $C(T)$-valued $X$ is also considered in [29]. There, $T=[0,t^{*}]$ is an interval and $Y$ is given by a partial observation of $X$ up to some earlier time $s\in[0,t^{*}]$, that is $Y:=X_{|[0,s]}$. Under an additional assumption, the author then proves that the conditional distribution $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a Gaussian measure on $C(T)$ and that the map $y\mapsto\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is even continuous if the set of probability measures on $C(T)$ is equipped with a metric describing the weak convergence of probability measures. In addition, it is shown that this additional assumption is also necessary for the existence of such continuous conditional probabilities. In fact it is worth to mention that both results are actually first shown for the general observational model (5) and then applied to the $C(T)$-case, but the motivation of [29] clearly focuses on the $C(T)$-case.
@@
If both $E$ and $F$ are Hilbert spaces, the conditioning problem has also been investigated by different techniques. For example, in [33] formulas for the mean and covariance of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ are established in the full infinite dimensional setting in terms of both a) so-called shorted operators and b) suitable sequences of oblique projections. Moreover, [26] shows, besides many other results, that some of the formulas derived in [33] can be significantly simplified. Independently, [31, 17] established simple formulas for the the mean and covariance, which essentially resemble (3) and (4) in an abstract sense, under the additional assumption that $\operatorname{cov}Y$ is injective.
The approximation in the $C(T)$-case with $S=T$ has also been considered in the literature. For example, convergence rates for the conditional mean functions, that is for $m_{X|Y_{n}=y_{n}}\to m_{X|Y=y}$, have been derived in [47] for specific Gaussian processes and the results of [15] can also be used to obtain such rates for random observational points and additional, vanishing observational noise. We refer to [24, Section 5.1] for a detailed discussion, where we note that [15] provides the better results compared to the source [39] mentioned in [24], since a) a clipping step for the mean is avoided and b) also stronger norms for the convergence are considered. In addition, convergence rates for Sobolev type covariance functions can be derived using the fill-distance and results from kernel interpolation [49]. For a detailed discussion, we again refer to [24] and the references mentioned therein. More recently, [28] established a uniform convergence result for the conditional mean and covariance functions under rather general assumptions on the underlying space $T$ and the prior kernel $k$. In addition, this paper provides a small list of earlier, less general convergence results in the $C(T)$-case. Besides the $C(T)$-case, however, we are not aware of any convergence results. We view this gap as substantial since conditioning on observations that are not direct partial observations of the paths of $X$ are interesting for various practical reasons, see e.g. the discussions in [45, 35].
@@
for almost all observations $y$ of $Y$, where $y_{n}:=(\langle y_{1}^{\prime},y\rangle_{F^{\prime},F},\dots,\langle y_{n}^{\prime},y\rangle_{F^{\prime},F}$. In addition, we show that the means and covariance operators also converge, see Theorem 3.4 for these three statements. Moreover, we prove that the means and covariances of $\mathrm{P}_{X|Y_{n}}(\,\cdot\,|y_{n})$ can be computed analogously to the GP4ML case (3) and (4), see Theorems 3.1 and 3.4. Furthermore, we show that the set of $y\in F$, for which the means converge, is a measurable, affine linear subspace of $F$ with full measure and this can be used to construct a particular version of $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$, see Theorem 3.6. As the title of this paper indicates, the major work horse in our proofs is a Banach space-valued martingale approach, which is coupled with the relationship between regular conditional probabilities and conditional expectations in the Banach space case, see Theorem 5.10. In this respect we note that martingale arguments have already been used by e.g. [31, 33] in the Hilbert space setting, but as far as we know, not in the general Banach space setting. In addition, the resulting martingale approximations in [31, 33] are restricted to rather specific filtrations, whereas our results only require a natural richness assumption on the filtration. Finally, we apply our general theory to the $C(T)$-valued case, where we establish uniformly convergent limit versions of the update rules (3) and (4) for general $S\subset T$.
The rest of this paper is organized as follows: In Section 2 we recall all notions and notations from probability theory and functional analysis that are necessary to understand our main results. These main results are then presented in Section 3 and the $C(T)$-valued case is subsequently investigated in Section 4. Sections 5 and 6 contain the proofs of the results of Section 3, respectively Section 4. Finally, we have attached four detailed appendices on A) Banach space valued integration, B) covariance operators, C) characteristic functions for Banach space valued random variables, and D) regular conditional probabilities and conditional expectations. Most material in these appendices is simply collected and unified from the existing literature, but a couple of mostly simple results are also added since we could not find them in the literature.

### v3

@@ Abstract
In this paper we investigate the conditional distributions of two Banach space valued, jointly Gaussian random variables. We show that these conditional distributions are again Gaussian and that their means and covariances are determined by a general finite dimensional approximation scheme based upon a martingale approach. In particular, it turns out that the covariance operators occurring in this scheme converge with respect to the nuclear norm and that the conditional probabilities converge weakly. Moreover, we discuss in detail, how our approximation scheme can be implemented in several classes of important Banach spaces such as (reproducing kernel) Hilbert spaces and spaces of continuous functions. As an example, we then apply our general results to the case of Gaussian processes with continuous paths conditioned to partial but infinite observations of their paths. Here we show that conditioning on sufficiently rich, increasing sets of finitely many observations leads to consistent approximations, that is, both the mean and covariance functions converge uniformly and the conditional probabilities converge weakly. Moreover, we discuss how these results improve our understanding of the popular Gaussian processes for machine learning.
Mathematical Subject Classification (2010). Primary 60G15; Secondary 60B10, 60B11, 62F15, 62G05, 62G20, 28C20, 68T05.
Key Words. Conditional Gaussian distributions in Banach spaces, Gaussian processes for machine learning, Probabilistic numerics
@@ 1 Introduction
Gaussian processes for machine learning (GP4ML), see e.g. [39, 7, 27], are a standard and widespread non-parametric Bayesian technique that allows for both prediction and uncertainty quantification in a single approach. In addition, their Bayesian update rule also plays an important role in other fields such as information-based complexity, see e.g. [48], probabilistic numerics and uncertainty quantification, see e.g. [46, 11, 23], Bayesian optimization, see e.g. [18], and Bayesian methods for inverse problems, see e.g. [45]. Finally, beginning with [35] they are also interesting for neural networks, where they are often named as neural network Gaussian processes (NNGPs).
In the simplest setup, GP4MLs assume that a Gaussian process $X:=(X_{t})_{t\in T}$ is given, where $T$ is some index set. Moreover, it is assumed that both the mean function $m:T\to\mathbb{R}$ and the covariance function $k:T\times T\to\mathbb{R}$ defined by
@@
for all $t,t_{1},t_{2}\in T$, where $K_{t_{1},S}$, $K_{S,S}$, and $K_{S,t_{2}}$ denote covariance matrices with respect to the indicated indices and $K_{S,S}^{\dagger}$ is the Moore-Penrose inverse of the $n\times n$-matrix $K_{S,S}$. Note that the GP4ML literature often assumes noisy observations of the form $Y+\varepsilon$, where $\varepsilon=(\varepsilon_{1},\dots,\varepsilon_{n})$ has i.i.d. entries with $\varepsilon_{i}\sim{\cal N}(0,\sigma^{2})$ and $\varepsilon$ is independent of $X$. In this case, $K_{S,S}^{\dagger}$ needs to be replaced by $(K_{S,S}+\sigma^{2}I_{n})^{-1}$, where $I_{n}$ denotes the $n$-dimensional identity matrix, see e.g. [39, Chapter 2.2] and [7, Chapter 6.4], as well as the discussion following (32).
@@
In the GP4ML literature, the Gaussian process $X$ is usually viewed as a mechanism to generate random functions, e.g. random continuous functions on a compact metric space $T$. In this case, the “conditional Gaussian process” of $X$ given the observations $Y=y$ is also viewed as a mechanism generating random continuous functions. To make this mathematically rigorous, one needs to view $X$ as a $C(T)$-valued Gaussian random variable, where $C(T)$ denotes the Banach space of continuous $\mathbb{R}$-valued functions on $T$, see Sections 2 and 5 for these and the following notions. Moreover, the above $Y$ is an $\mathbb{R}^{n}$-valued Gaussian random variable such that $X$ and $Y$ are jointly Gaussian. Given their joint distribution $\mathrm{P}_{(X,Y)}$ on $C(T)\times\mathbb{R}^{n}$, one then seeks a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$, since $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is the conditional distribution of $X$ given the observation $Y=y$. Note that by definition, each $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a distribution on $C(T)$ and therefore we can also ensure that the “conditional Gaussian process” again produces continuous functions.
@@
The references mentioned at the beginning indicate that for the conditioning problem $C(T)$ is by no means the only Banach space one is interested in. Therefore, let us now consider the general case. To this end, let $E$ and $F$ be Banach spaces, which due to technical simplicity are assumed to be separable. Moreover, let $X$ be an $E$-valued random variable and $Y$ be an $F$-valued random variable such that $X$ and $Y$ are jointly Gaussian. In the abstract setting one then seeks a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$. Again, $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a distribution on $E$, which equals the conditional distribution of $X$ given the observation $Y=y$. Compared to the $C(T)$-case described earlier, this generalization opens several additional possibilities. First, of course, we can consider Banach spaces of functions $E$ other than $C(T)$, e.g. spaces of differentiable functions. Second, we no longer need to assume that we can observe the random functions on some finite number of observational points. Instead, we could assume, e.g., that we observe derivatives of the random functions at some observational points, cf. [38]. More generally, $Y$ can, for example, be composed by any other finite set of bounded linear functionals as considered in e.g. [49], or even by an infinite set of such functionals. In particular, for the $C(T)$-example above, the abstract setting makes it possible to consider arbitrary subsets $S\subset T$ for the observational points, which is important for ODE- and PDE-related applications, see e.g. [23, 38] and the references mentioned therein. To abstract from these examples, let us assume that $Y$ is given by the “observational model”
@@
Although working with infinite dimensional $Y$ has clear conceptual advantages, there is also a drawback: Indeed, in many situations, it is practically impossible to fully observe or process infinite dimensional $Y$, or to even compute $\mathrm{P}_{X|Y}$ for such $Y$. In these cases, one often uses a sequence $(Y_{n})$ of observable, finite dimensional Gaussian random variables, for which the condition step is feasible. For example, in the $C(T)$-case with infinite $S\subset T$, one could choose an increasing sequence $(S_{n})$ of finite subsets $S_{n}\subset S$, whose union is dense in $S$, and set $Y_{n}:=(X_{s})_{s\in S_{n}}$. The hope of this approach is, of course, that $\mathrm{P}_{X|Y_{n}}$ provides a convergent approximation of $\mathrm{P}_{X|Y}$, e.g. in terms of their means and covariances, or as distributions.
@@
Both the fully infinite dimensional conditioning and its possible approximations by finite dimensional observations have various applications as described in e.g. [48, 11, 23, 38] and the references mentioned therein. Here, we only mention Bayesian quadrature, Bayesian optimization, and Bayesian solvers for differential equations. In addition, they can help to better understand GP4MLs and their limiting behavior for $n\to\infty$. Let us therefore have a look what is known so far.
@@
To begin with, in the case of finite dimensional observations, that is $F=\mathbb{R}^{n}$, [8, Corollary 3.10.3] shows that there is a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$, such that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian for all $y$. However, no formulas for the mean and covariance of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ are provided. Moreover, the observational model (5) for centered $X$ and with arbitrary $E$ and $F$ has been investigated in [47], where the authors show that the regular conditional probabilities are Gaussian measures, see their Theorem 3.11. Furthermore, in the case $F=\mathbb{R}^{n}$, their Corollary 3.13 present explicit formulas for the conditional mean and covariance with the help of a so-called representing sequence of the covariance operator $\operatorname{cov}Y$. In addition, a closer look at their proofs reveals, that analogous formulas hold true for general, separable $F$. The latter formulas have been made explicit and generalized to non-centered $X$ in [49]. Moreover, both papers establish the linearity of the conditional mean mapping, and [49] also describes formulas for sequential updates and the case of $C(T)$-valued $X$. In particular, for $E=C(T)$ and $F=\mathbb{R}^{n}$, the update rules (3) and (4) are generalized to arbitrary, bounded linear observational operators $\mathsf{S}:C(T)\to\mathbb{R}^{n}$. Unfortunately, however, the mean and covariance formulas for infinite dimensional $F$ become less handy, as they require to have a representing sequence for $\operatorname{cov}Y$. Although such sequences as well as constructions for them exist, see e.g. [47, Lemma 3.5] or [44, Lemma 8.2.3], they are rarely at hand in situations in which $F$ is not a Hilbert space. As a consequence, it seems fair to say that the derived formulas have some limited practical value.
@@
The case of $C(T)$-valued $X$ is also considered in [31]. There, $T=[0,t^{*}]$ is an interval and $Y$ is given by a partial observation of $X$ up to some earlier time $s\in[0,t^{*}]$, that is $Y:=X_{|[0,s]}$. Under an additional assumption, the author then proves that the conditional distribution $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a Gaussian measure on $C(T)$ and that the map $y\mapsto\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is even continuous if the set of probability measures on $C(T)$ is equipped with the topology describing the weak convergence of probability measures. Furthermore, it is shown that this additional assumption is also necessary for the existence of such continuous conditional probabilities. In fact it is worth to mention that both results are actually first shown for the general observational model (5) and then applied to the $C(T)$-case, but the motivation of [31] clearly focuses on the $C(T)$-case.
@@
If both $E$ and $F$ are Hilbert spaces, the conditioning problem has also been investigated by different techniques. For example, in [36] formulas for the mean and covariance of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ are established in the full infinite dimensional setting in terms of a) so-called shorted operators, and b) suitable sequences of oblique, i.e. non-orthogonal, projections. Moreover, [28] shows, besides many other results, that some of the formulas derived in [36] can be significantly simplified. Independently, [33, 20] established simple formulas for the the mean and covariance, which essentially resemble (3) and (4) in an abstract sense, under the additional assumption that $\operatorname{cov}Y$ is injective.
The approximation in the $C(T)$-case with $S=T$ has also been considered in the literature. For example, convergence rates for the conditional mean functions, that is for $m_{X|Y_{n}=y_{n}}\to m_{X|Y=y}$, have been derived in [51] for specific Gaussian processes and the results of [17] can also be used to obtain such rates for random observational points and additional, vanishing observational noise. We refer to [27, Section 5.1] for a detailed discussion, where we note that [17] provides the better results compared to the source [43] mentioned in [27], since a) a clipping step for the mean is avoided, and b) also stronger norms for the convergence are considered. In addition, convergence rates for Sobolev type covariance functions can be derived using the fill-distance and results from kernel interpolation [53]. For a detailed discussion, we again refer to [27] and the references mentioned therein. More recently, [30] established a uniform convergence result for the conditional mean and covariance functions under rather general assumptions on the underlying space $T$ and the prior kernel $k$. In addition, this paper provides a small list of earlier, less general convergence results in the $C(T)$-case with $S=T$. Besides this case, however, there are, to the best of our knowledge, no other established convergence results. This stands in contrast to the variety of possible applications mentioned above, see [48, 47, 46, 23], as well as in e.g. [49, 38].
@@
for almost all observations $y$ of $Y$, where $y_{n}:=(\langle y_{1}^{\prime},y\rangle_{F^{\prime},F},\dots,\langle y_{n}^{\prime},y\rangle_{F^{\prime},F}$. In addition, we show that the mean vectors and covariance operators also converge, see Theorem 3.3 for these three statements. Moreover, we prove that the means and covariances of $\mathrm{P}_{X|Y_{n}}(\,\cdot\,|y_{n})$ can be computed analogously to the GP4ML case (3) and (4), see Theorems 3.1 and 3.3. Furthermore, we show that the set of $y\in F$, for which the means converge, is a measurable, affine linear subspace of $F$ with full measure and this can be used to construct a particularly “benign” version of $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$, see Theorem 3.5. As the title of this paper indicates, the major work horse in our proofs is a Banach space-valued martingale approach, which is coupled with the relationship between regular conditional probabilities and conditional expectations in the Banach space case, see Theorem 6.7. In this respect we note that martingale arguments have already been used by e.g. [33, 36] in the Hilbert space setting, but as far as we know, not in the general Banach space setting. In addition, the resulting martingale approximations in [33, 36] are restricted to rather specific filtrations, whereas our results only require a natural richness assumption on the filtration. Finally, we apply our general theory to the $C(T)$-valued case, where we establish uniformly convergent limit versions of the update rules (3) and (4) for general $S\subset T$. These results generalize the findings of [30] in the sense that both $S=T$ is no longer required and no additional assumptions on the observational model are needed.
The rest of this paper is organized as follows: In Section 2 we recall all notions and notations from probability theory and functional analysis that are necessary to understand our main results. These main results are then presented in Section 3. Moreover, in Section 4 we investigate the required “sufficiently rich” sequence $(y_{n}^{\prime})\subset F^{\prime}$ mentioned above in more detail. In particular, we establish a general existence result and discuss simple constructions for some concrete types of spaces. The $C(T)$-valued case is subsequently investigated in Section 5. Sections 6 to 8 contain the proofs of the results of Sections 3 to 5. Finally, we have attached four detailed appendices on A) Banach space valued integration, B) covariance operators, C) characteristic functions for Banach space valued random variables, and D) regular conditional probabilities and conditional expectations in the Banach space case. Most material in these appendices is simply collected and unified from the existing literature, but a couple of mostly simple results are also added since we could not find them in the literature.